OpenStudy (anonymous):

Someone please help Verify the identity cot ( θ - pi/2) = - tan θ

OpenStudy (anonymous):

ok i will...:)

OpenStudy (anonymous):

what is cot(x-y)?...cotxcoty+1/coty-cotx...right?

OpenStudy (anonymous):

put x=theta and y=pi/2

OpenStudy (anonymous):

we get \[\cot(\theta-\pi/2)=\cot \theta.(\cot \pi/2)/(\cot \pi/2)-\cot \theta\]

OpenStudy (anonymous):

cot pi/2=0..

OpenStudy (anonymous):

we get the answer as 1/-cot theta which is -tan theta

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

My pleasure @Aurelia0912